Consider a system with a damping force undergoing forced oscillations at an angular frequency omega

Find the instantaneous kinetic and potential energy of the system.


1
Expert's answer
2021-02-22T10:21:46-0500

Answer

Displacement of damped harmonic oscillator is given

"x=ae^{-bt}sin(wt+\phi)"

Where

"w=\sqrt{w_0^2-b^2}"

Now kinetic energy is given by

"K=\frac{m}{2}(\frac{dx}{dt}) ^2"

Putting displacement x and differentiate with respect to t

"K=\frac{ma^2e^{-2bt}}{2}(b^2sin^2(wt+\phi) +w^2cos^2(wt+\phi) -2bwsin(wt+\phi) cos(wt+\phi))"


Now potential energy

"P=\frac{mw_0^2x^2}{2}"

Now putting value of displacement

"P=\frac{a^2mw_0^2e^{-2bt}sin^2(wt+\phi) }{2}"